# Repeating Flares, X-ray Outbursts and Delayed Infrared Emission: A Comprehensive Compilation of Optical Tidal Disruption Events

### v1

@@ Acknowledgements.
- Sun et al. (2024) Sun, L., Jiang, N., Dou, L., et al. 2024, Recurring tidal disruption events a decade apart in IRAS F01004-2237
- Syer & Ulmer (1999) Syer, D. & Ulmer, A. 1999, MNRAS, 306, 35
- Tadhunter et al. (2017) Tadhunter, C., Spence, R., Rose, M., Mullaney, J., & Crowther, P. 2017, Nature Astronomy, 1, 0061
- Tonry et al. (2018) Tonry, J. L., Denneau, L., Heinze, A. N., et al. 2018, PASP, 130, 064505
- Tozzi et al. (2006) Tozzi, P., Gilli, R., Mainieri, V., et al. 2006, A&A, 451, 457
- van Velzen et al. (2016) van Velzen, S., Anderson, G. E., Stone, N. C., et al. 2016, Science, 351, 62
- van Velzen et al. (2021) van Velzen, S., Gezari, S., Hammerstein, E., et al. 2021, ApJ, 908, 4
- Virtanen et al. (2020) Virtanen, P., Gommers, R., Oliphant, T. E., et al. 2020, Nature Methods, 17, 261
- Wang et al. (2022) Wang, Y., Jiang, N., Wang, T., et al. 2022, The Astrophysical Journal Letters, 930, L4
- Waters et al. (2020) Waters, C. Z., Magnier, E. A., Price, P. A., et al. 2020, ApJS, 251, 4
- Webb et al. (2020) Webb, N. A., Coriat, M., Traulsen, I., et al. 2020, A&A, 641, A136
- Wevers et al. (2023) Wevers, T., Coughlin, E. R., Pasham, D. R., et al. 2023, ApJ, 942, L33
- Wevers et al. (2019) Wevers, T., Pasham, D. R., van Velzen, S., et al. 2019, MNRAS, 488, 4816
- Wright et al. (2010) Wright, E. L., Eisenhardt, P. R. M., Mainzer, A. K., et al. 2010, AJ, 140, 1868
- Wyrzykowski et al. (2017) Wyrzykowski, Ł., Zieliński, M., Kostrzewa-Rutkowska, Z., et al. 2017, MNRAS, 465, L114
- Yao & Guolo (2022) Yao, Y. & Guolo, M. 2022, The Astronomer’s Telegram, 15657, 1
- Yao et al. (2023) Yao, Y., Ravi, V., Gezari, S., et al. 2023, ApJ, 955, L6
- Yaron & Gal-Yam (2012) Yaron, O. & Gal-Yam, A. 2012, PASP, 124, 668
- York et al. (2000) York, D. G., Adelman, J., Anderson, Jr., J. E., et al. 2000, AJ, 120, 1579
- Yuan et al. (2024) Yuan, C., Winter, W., & Lunardini, C. 2024, ApJ, 969, 136
- Zauderer et al. (2011) Zauderer, B. A., Berger, E., Soderberg, A. M., et al. 2011, Nature, 476, 425
- Zhong (2025) Zhong, S. 2025, ApJ, 983, 131
@@ Appendix A Published TDEs
AT 2017gge: This TDE was first reported in Wang et al. (2022) (ATLAS17jrp) and later appeared in Onori et al. (2022). Recently, it was reported to be the second TDE associated with a high-energy neutrino in Li et al. (2024), following AT 2019dsg (Stein et al. 2021b).
@@
PTF09djl, PTF09ge: These TDEs were first presented in Arcavi et al. (2014), where they were characterized as TDE candidates. Later, they were included in the TDE sample of Charalampopoulos et al. (2022).
@@
ASASSN-14li: This TDE was discovered in December of 2014 (Jose et al. 2014) and has since been studied extensively across the electromagnetic spectrum, in X-rays (Miller et al. 2015; Holoien et al. 2016b; Pasham et al. 2017), in the optical and near-ultraviolet (UV) (Cenko et al. 2016; Holoien et al. 2016b; Pasham et al. 2017), infrared (IR) (Jiang et al. 2016) and radio (Alexander et al. 2016; van Velzen et al. 2016). It was later included in the Charalampopoulos et al. (2022) TDE sample.
@@ B.1.1 Data Reduction
The XRT photon counting (PC) mode data were downloaded from HEASARC161616https://heasarc.gsfc.nasa.gov/ data archive, and processed using the XRTDAS software (Capalbi et al. 2005) developed at the ASI Science Data Center and included in the HEAsoft package (v. 6.33.2) distributed by HEASARC, using a procedure similar to that illustrated in Paggi et al. (2013).
For each observation calibrated and cleaned PC mode event files were produced with the xrtpipeline task (ver. 0.13.7), also producing exposure maps for each observation. In addition to the screening criteria used by the standard pipeline processing, we applied a further filter to screen background spikes that can occur when the angle between the pointing direction of the satellite and the bright Earth limb is low. In order to eliminate this so called bright Earth effect, due to the scattered optical light that usually occurs towards the beginning or the end of each orbit, we used the procedure proposed by Puccetti et al. (2011) and D’Elia et al. (2013). We monitored the count-rate on the CCD border and, through the xselect package, we excluded time intervals when the count-rate in this region exceeded . In addition we selected only time intervals with CCD temperatures less than (instead of the standard limit of ) since contamination by dark current and hot pixels, which increase the low energy background, is strongly temperature dependent (D’Elia et al. 2013).
@@ B.1.2 Source Detection and Flux estimates
To detect X-ray sources in the XRT images, we made use of the ximage detection algorithm detect, which locates the point sources using a sliding-cell method. The average background intensity is estimated in several small square boxes uniformly located within the image. The position and intensity of each detected source are calculated in a box whose size maximizes the signal-to-noise ratio.
For each XRT-PC observation, we considered as coordinates of the X-ray counterpart to the TDE source the coordinates of the detected XRT-PC source closest to the TDE source coordinates if this happens to lie closer than . If no XRT-PC source was detected closer than to the TDE source coordinates, we used the TDE source coordinates themselves.
We then evaluated net count-rates (or their upper limits) at X-ray counterpart coordinates with the sosta algorithm that, besides the net count-rates and the respective uncertainties, yields the statistical significance of each source. In addition, sosta also estimates the optimal extraction radius that maximizes the signal to noise ratio of the source. We note that we used count-rates produced by sosta rather than those given by detect because the former are in most cases more accurate, since detect uses a global background for the entire image, whereas sosta uses a local background.
In order to get a first estimate of fluxes for TDE sources X-ray counterparts, we extracted appropriate arf and rmf files at each source location, making use of the xrtproducts task. As extraction regions we used circles centred at the X-ray source coordinates and with a radius equal to . Assuming black body model with a temperature of and an absorption component fixed to the Galactic value (Kalberla et al. 2005), we then converted the net count-rates evaluated earlier in intrinsic (i.e. unabsorbed) fluxes.
@@ B.1.3 Spectral Extraction
In general, source spectra - with the corresponding arf and rmf files - were extracted form events with xrtproducts task, using circular regions centred at the X-ray source coordinates with radii equal , while background spectra were extracted from annuli centred at the X-ray source coordinates, with inner and outer radii equal to and , respectively, excluding nearby detected sources.
When the source count-rate is above , the data are significantly affected by pileup in the inner part of the PSF (Moretti et al. 2005). To remove the pile-up contamination, we extract only events contained in an annular region centred on the X-ray source coordinates (Perri et al. 2007). While the outer radius of the annulus was set at , the inner radius was determined by comparing the observed profiles with the updated XRT PSF analytical model171717https://www.swift.ac.uk/analysis/xrt/SWIFT-XRT-CALDB-10_v01.pdf.
@@ B.2.1 Data Reduction
Chandra-ACIS data were retrieved from the Chandra Data Archive181818http://cda.harvard.edu/chaser, and were processed and analysed with the Chandra Interactive Analysis of Observations (CIAO, Fruscione et al. 2006) data analysis system version 4.16 and Chandra calibration database CALDB version 4.11.1, adopting standard procedures. We run the ACIS level 2 processing with chandra_repro to apply up-to-date calibrations (CTI correction, ACIS gain, bad pixels), and then excluded time intervals of background flares exceeding with the deflare task. We produced full-band exposure maps, psf maps, and pileup maps with the pileup_map task.
@@ B.2.2 Source Detection and Flux estimates
To evaluate the source fluxes we made use of the srcflux task, extracting counts from circular regions centered at the source location, with radii equal to the 99% EEF PSF radius (as estimated from the PSF maps), while for the background regions we considered annuli with inner and outer radii equal to and , respectively,excluding nearby detected sources. Appropriate arf and rmf files were generated for the considered regions, and the net count-rates were converted to intrinsic fluxes assuming a model comprising an absorption component fixed to the Galactic value and a black body with a temperature of . In the case of unconstrained count-rates, upper limits were evaluated with the aprates_tool task, and converted to intrinsic flux upper limits with the modelflux task, adopting the same model described before.
@@ B.2.3 Spectral Extraction
Source spectra and the corresponding arf and rmf files were extracted form event files with specextract task, using circular regions centred at the X-ray source coordinates with radii equal , while background spectra were extracted from annuli centred at the X-ray source coordinates, with inner and outer radii equal to and , respectively, excluding nearby detected sources. Also in this case we excluded inner pixels with pileup larger than as estimated from the pileup maps.
@@ B.3.1 Data Reduction
XMM-Newton-EPIC data were retrieved from the XMM-Newton Science Archive191919http://nxsa.esac.esa.int/nxsa-web and reduced with the SAS202020http://www.cosmos.esa.int/web/xmm-newton/sas 21.0.0 software.
Following Nevalainen et al. (2005) we filtered EPIC data for hard-band flares by excluding the time intervals where the (for MOS1 and MOS2) or (for PN) count-rate evaluated on the whole detector FOV was more than 3 away from its average value. To achieve a tighter filtering of background flares, we iteratively repeated this process two more times, re-evaluating the average hard-band count-rate and excluding time intervals away more than from this value. The same procedure was applied to soft band restricting the analysis to an annulus with inner and outer radii of and excluding sources in the field, where the detected emission is expected to be dominated by the background.
@@ B.3.2 Source Detection and Flux estimates
When possible, we merged data from MOS1, MOS2 and PN detectors using the emosaic task, in order to detect the fainter sources that wouldn’t be detected otherwise. Sources were detected following the standard SAS sliding box task edetect_chain that mainly consist of three steps: 1) source detection with local background, with a minimum detection likelihood of 8; 2) remove sources in step 1 and create a smooth background maps by fitting a 2-D spline to the residual image; 3) source detection with the background map produced in step 2 with a minimum detection likelihood of 10. The task emldetect was then used to determine the parameters for each input source - including the count-rate - by means of a maximum likelihood fit to the input images, selecting sources with a minimum detection likelihood of 15 and a flux in the band larger than (assuming an energy conversion factor of ). An analytical model of the PSF was evaluated at the source position and normalized to the source brightness. The source extent was then evaluated as the radius at which the PSF level equals half of local background.
Again, we considered as coordinates of the X-ray counterpart to the TDE source the coordinates of the detected XMM-Newton-EPIC source closest to the TDE source coordinates if this happens to lie closer than . If no XMM-Newton-EPIC source was detected closer than to the TDE source coordinates, we considered as coordinates of the possible X-ray counterpart the TDE source coordinates themselves. In this case, the source net count-rate (or its upper limits) and optimal extraction radius (maximizing the signal to noise ratio of the source) was provided by eregionanalyse.
Again, to estimate the fluxes of TDE sources X-ray counterparts, we extracted appropriate arf and rmf files for each available detector at each source location, making use of the arfgen and rmfgen tasks. As extraction regions we used circles centred at the X-ray source coordinates and with a radius equal to . Assuming a black body model with a temperature of and a absorption component fixed to the Galactic value, we then converted the net count-rates of each detector intrinsic (i.e. unabsorbed) fluxes. The mean source flux was then finally evaluated as the mean of the fluxes of each available detector weighted by its uncertainty, and the uncertainty on this mean flux was evaluated as:

### v2

@@ Acknowledgements.
- Recurring tidal disruption events a decade apart in iras f01004-2237. External Links: 2410.09720, Link Cited by: Appendix A, 2nd item, §4.3, §4.3, footnote 16.
- Tidal disruption rates of stars in observed galaxies. MNRAS 306 (1), pp. 35–42. External Links: Document, astro-ph/9812389, ADS entry Cited by: §5.2.
- A tidal disruption event in the nearby ultra-luminous infrared galaxy F01004-2237. Nature Astronomy 1, pp. 0061. External Links: Document, 1702.02573, ADS entry Cited by: Appendix A.
- ATLAS: A High-cadence All-sky Survey System. PASP 130 (988), pp. 064505. External Links: Document, 1802.00879, ADS entry Cited by: §1.
- A radio jet from the optical and x-ray bright stellar tidal disruption flare ASASSN-14li. Science 351 (6268), pp. 62–65. External Links: Document, 1511.08803, ADS entry Cited by: Appendix A.
- Optical Discovery of Probable Stellar Tidal Disruption Flares. ApJ 741 (2), pp. 73. External Links: Document, 1009.1627, ADS entry Cited by: §1.
- Seventeen Tidal Disruption Events from the First Half of ZTF Survey Observations: Entering a New Era of Population Studies. ApJ 908 (1), pp. 4. External Links: Document, 2001.01409, ADS entry Cited by: §1, §2.1, §2.1, Table 1, §3.4, §4.2, §4.2, §5.1, §5.2, §5.3.
- SciPy 1.0: fundamental algorithms for scientific computing in Python. Nature Methods 17, pp. 261–272. External Links: Document, 1907.10121, ADS entry Cited by: Data availability.
- Revised Rates of Stellar Disruption in Galactic Nuclei. ApJ 600 (1), pp. 149–161. External Links: Document, astro-ph/0305493, ADS entry Cited by: §1.
- Discovery of atlas17jrp as an optical-, x-ray-, and infrared-bright tidal disruption event in a star-forming galaxy. The Astrophysical Journal Letters 930 (1), pp. L4. External Links: Document, Link Cited by: Appendix A, 3rd item.
- Pan-STARRS Pixel Processing: Detrending, Warping, Stacking. ApJS 251 (1), pp. 4. External Links: Document, 1612.05245, ADS entry Cited by: §1.
- The XMM-Newton serendipitous survey. IX. The fourth XMM-Newton serendipitous source catalogue. A&A 641, pp. A136. External Links: Document, 2007.02899, ADS entry Cited by: §B.3.2, §B.5.1, §3.2.
- Live to Die Another Day: The Rebrightening of AT 2018fyk as a Repeating Partial Tidal Disruption Event. ApJ 942 (2), pp. L33. External Links: Document, 2209.07538, ADS entry Cited by: footnote 16.
- An elliptical accretion disk following the tidal disruption event AT 2020zso. A&A 666, pp. A6. External Links: Document, 2202.08268, ADS entry Cited by: §1.
- Evidence for rapid disc formation and reprocessing in the X-ray bright tidal disruption event candidate AT 2018fyk. MNRAS 488 (4), pp. 4816–4830. External Links: Document, 1903.12203, ADS entry Cited by: footnote 16.
- The Wide-field Infrared Survey Explorer (WISE): Mission Description and Initial On-orbit Performance. AJ 140 (6), pp. 1868–1881. External Links: Document, 1008.0031, ADS entry Cited by: §1.
- OGLE16aaa - a signature of a hungry supermassive black hole. MNRAS 465 (1), pp. L114–L118. External Links: Document, 1606.03125, ADS entry Cited by: Table 1.
- Tidal Disruption Event Demographics with the Zwicky Transient Facility: Volumetric Rates, Luminosity Function, and Implications for the Local Black Hole Mass Function. ApJ 955 (1), pp. L6. External Links: Document, 2303.06523, ADS entry Cited by: §1, §2.1, §2.1, §4.2, §4.3, §5.1, §5.3, footnote 16.
- WISeREP—An Interactive Supernova Data Repository. PASP 124 (917), pp. 668. External Links: Document, 1204.1891, ADS entry Cited by: §3.4.
- The Sloan Digital Sky Survey: Technical Summary. AJ 120 (3), pp. 1579–1587. External Links: Document, astro-ph/0006396, ADS entry Cited by: §1.
- AT2021lwx: Another Neutrino-coincident Tidal Disruption Event with a Strong Dust Echo?. ApJ 969 (2), pp. 136. External Links: Document, 2401.09320, ADS entry Cited by: §5.4.
- Birth of a relativistic outflow in the unusual -ray transient Swift J164449.3+573451. Nature 476 (7361), pp. 425–428. External Links: Document, 1106.3568, ADS entry Cited by: Appendix A.
- Modeling the UV/Optical Light Curve of Re-brightening Tidal Disruption Events. ApJ 983 (2), pp. 131. External Links: Document, 2412.12549, ADS entry Cited by: §5.3, footnote 16.
@@ Appendix A Published TDEs
AT 2017gge: This TDE was first reported in Wang et al. (2022) (ATLAS17jrp) and later appeared in Onori et al. (2022). Recently, it was reported to be the second TDE associated with a high-energy neutrino in Li et al. (2024), following AT 2019dsg (Stein et al. 2021).
@@
PTF09djl, PTF09ge: These TDEs were first presented in Arcavi et al. (2014), where they were characterised as TDE candidates. Later, they were included in the TDE sample of Charalampopoulos et al. (2022).
@@
ASASSN-14li: This TDE was discovered in December of 2014 (Jose et al. 2014) and has since been studied extensively across the electromagnetic spectrum, in X-rays (Miller et al. 2015; Holoien et al. 2016b; Pasham et al. 2017), in the optical and near-UV (Cenko et al. 2016; Holoien et al. 2016b; Pasham et al. 2017), IR (Jiang et al. 2016) and radio (Alexander et al. 2016; van Velzen et al. 2016). It was later included in the Charalampopoulos et al. (2022) TDE sample.
@@ B.1.1 Data reduction
The XRT photon counting (PC) mode data were downloaded from HEASARC212121https://heasarc.gsfc.nasa.gov/ data archive, and processed using the XRTDAS software (Capalbi et al. 2005) developed at the ASI Science Data Center and included in the HEAsoft package (v. 6.33.2) distributed by HEASARC, using a procedure similar to that illustrated in Paggi et al. (2013).
For each observation calibrated and cleaned PC mode event files were produced with the xrtpipeline task (ver. 0.13.7), also producing exposure maps for each observation. In addition to the screening criteria used by the standard pipeline processing, we applied a further filter to screen background spikes that can occur when the angle between the pointing direction of the satellite and the bright Earth limb is low. In order to eliminate this so called bright Earth effect, due to the scattered optical light that usually occurs towards the beginning or the end of each orbit, we used the procedure proposed by Puccetti et al. (2011) and D’Elia et al. (2013). We monitored the count rate on the CCD border and, through the xselect package, we excluded time intervals when the count rate in this region exceeded . In addition we selected only time intervals with CCD temperatures less than (instead of the standard limit of ) since contamination by dark current and hot pixels, which increase the low energy background, is strongly temperature dependent (D’Elia et al. 2013).
@@ B.1.2 Source detection and flux estimates
To detect X-ray sources in the XRT images, we made use of the ximage detection algorithm detect, which locates the point sources using a sliding-cell method. The average background intensity is estimated in several small square boxes uniformly located within the image. The position and intensity of each detected source are calculated in a box whose size maximises the S/N.
For each XRT-PC observation, we considered as coordinates of the X-ray counterpart to the TDE source the coordinates of the detected XRT-PC source closest to the TDE source coordinates if this happens to lie closer than 222222This radius represents a conservative value that includes systematic astrometric uncertainties of the various telescopes, as well as positional uncertainties related to the source fluxes. We checked that the majority of the detected X-ray sources fall at separations well below this value, with 70% of the sources at separations , 85% of the sources at separations , and more than 90% of the sources at separations .. If no XRT-PC source was detected closer than to the TDE source coordinates, we used the TDE source coordinates themselves.
We then evaluated net count rates (or their upper limits) at X-ray counterpart coordinates with the sosta algorithm that, besides the net count rates and the respective uncertainties, yields the statistical significance of each source. In addition, sosta also estimates the optimal extraction radius that maximises the S/N of the source. We note that we used count rates produced by sosta rather than those given by detect because the former are in most cases more accurate, since detect uses a global background for the entire image, whereas sosta uses a local background.
In order to get a first estimate of fluxes for TDE sources X-ray counterparts, we extracted appropriate arf and rmf files at each source location, making use of the xrtproducts task. As extraction regions we used circles centred at the X-ray source coordinates and with a radius equal to . Assuming black body model with a temperature of K232323This value has been evaluated from the error weighted average temperature obtained from the spectral fits, see Appendix B.4 and an absorption component fixed to the Galactic value (Kalberla et al. 2005), we then converted the net 0.3 - 10 keV count rates evaluated earlier in 0.3 - 10 keV observed intrinsic (i.e. unabsorbed) fluxes.
@@ B.1.3 Spectral extraction
In general, source spectra - with the corresponding arf and rmf files - were extracted form events with xrtproducts task, using circular regions centred at the X-ray source coordinates with radii equal , while background spectra were extracted from annuli centred at the X-ray source coordinates, with inner and outer radii equal to and , respectively, excluding nearby detected X-ray sources.
When the source count rate is above , the data are significantly affected by pileup in the inner part of the PSF (Moretti et al. 2005). To remove the pile-up contamination, we extract only events contained in an annular region centred on the X-ray source coordinates (Perri et al. 2007). While the outer radius of the annulus was set at , the inner radius was determined by comparing the observed profiles with the updated XRT PSF analytical model242424https://www.swift.ac.uk/analysis/xrt/SWIFT-XRT-CALDB-10_v01.pdf.
@@ B.2.1 Data reduction
Chandra-ACIS data were retrieved from the Chandra Data Archive252525http://cda.harvard.edu/chaser, and were processed and analysed with the Chandra Interactive Analysis of Observations (CIAO, Fruscione et al. 2006) data analysis system version 4.16 and Chandra calibration database CALDB version 4.11.1, adopting standard procedures. We run the ACIS level 2 processing with chandra_repro to apply up-to-date calibrations (CTI correction, ACIS gain, bad pixels), and then excluded time intervals of background flares exceeding with the deflare task. We produced full-band exposure maps, psf maps, and pileup maps with the pileup_map task.
@@ B.2.2 Source detection and flux estimates
To evaluate the source fluxes we made use of the srcflux task, extracting counts from circular regions centered at the source location, with radii equal to the 99% EEF PSF radius (as estimated from the PSF maps), while for the background regions we considered annuli with inner and outer radii equal to and , respectively, excluding nearby detected X-ray sources. Appropriate arf and rmf files were generated for the considered regions, and the net count rates were converted to observed intrinsic fluxes assuming a model comprising an absorption component fixed to the Galactic value and a black body with a temperature of . In the case of unconstrained count rates, upper limits were evaluated with the aprates_tool task, and converted to intrinsic flux upper limits with the modelflux task, adopting the same model described before.
@@ B.2.3 Spectral extraction
Source spectra and the corresponding arf and rmf files were extracted form event files with specextract task, using circular regions centred at the X-ray source coordinates with radii equal , while background spectra were extracted from annuli centred at the X-ray source coordinates, with inner and outer radii equal to and , respectively, excluding nearby detected X-ray sources. Also in this case we excluded inner pixels with pileup larger than as estimated from the pileup maps.
@@ B.3.1 Data reduction
XMM-Newton-EPIC data were retrieved from the XMM-Newton Science Archive262626http://nxsa.esac.esa.int/nxsa-web and reduced with the SAS272727http://www.cosmos.esa.int/web/xmm-newton/sas 21.0.0 software.
Following Nevalainen et al. (2005) we filtered EPIC data for hard-band flares by excluding the time intervals where the (for MOS1 and MOS2) or (for PN) count rate evaluated on the whole detector FOV was more than 3 away from its average value. To achieve a tighter filtering of background flares, we iteratively repeated this process two more times, re-evaluating the average hard-band count rate and excluding time intervals away more than from this value. The same procedure was applied to soft band restricting the analysis to an annulus with inner and outer radii of and excluding detected X-ray sources in the field, where the detected emission is expected to be dominated by the background.
@@ B.3.2 Source detection and flux estimates
When possible, we merged data from MOS1, MOS2 and PN detectors using the emosaic task, in order to detect the fainter sources that would not be detected otherwise. Sources were detected following the standard SAS sliding box task edetect_chain that mainly consist of three steps: 1) source detection with local background, with a minimum detection likelihood of 8; 2) remove sources in step 1 and create a smooth background maps by fitting a 2-D spline to the residual image; 3) source detection with the background map produced in step 2 with a minimum detection likelihood of 10. The task emldetect was then used to determine the parameters for each input source - including the count rate - by means of a maximum likelihood fit to the input images, selecting sources with a minimum detection likelihood of 15 and a flux in the band larger than (assuming an energy conversion factor of ). An analytical model of the PSF was evaluated at the source position and normalised to the source brightness. The source extent was then evaluated as the radius at which the PSF level equals half of local background.
Again, we considered as coordinates of the X-ray counterpart to the TDE source the coordinates of the detected XMM-Newton-EPIC source closest to the TDE source coordinates if this happens to lie closer than . If no XMM-Newton-EPIC source was detected closer than to the TDE source coordinates, we considered as coordinates of the possible X-ray counterpart the TDE source coordinates themselves. In this case, the source net count rate (or its upper limits) and optimal extraction radius (maximising the signal to noise ratio of the source) was provided by eregionanalyse.
Again, to estimate the fluxes of TDE sources X-ray counterparts, we extracted appropriate arf and rmf files for each available detector at each source location, making use of the arfgen and rmfgen tasks. As extraction regions we used circles centred at the X-ray source coordinates and with a radius equal to . Assuming a black body model with a temperature of and a absorption component fixed to the Galactic value, we then converted the net count rates of each detector intrinsic (i.e. unabsorbed) fluxes. The mean source flux was then finally evaluated as the mean of the fluxes of each available detector weighted by its uncertainty, and the uncertainty on this mean flux was evaluated as